OpenStudy (anonymous):

Solve by graphing y=1/2x+2 Y=x+1

OpenStudy (johnweldon1993):

So we're seeing where they intersect I suppose Instead of graphing you can just set the 2 equations = to each other y = 1/2x + 2 and y = x+1 so 1/2x + 2 = x + 1 get like terms on the same side 1/2x - x = 1-2 -1/2x = -1 divide both sides by -1/2 and you get x = 2 just plug 2 into either equation to find y which is 3

OpenStudy (anonymous):